# Online Smoothed Demand Management

### v1

@@ Appendix D Deferred Proofs from § 3
In the above, $\alpha_{\texttt{RORO}}$ is the optimal competitive ratio for OCS, which is defined in the context of OSDM as the unique positive solution to $\exp\left(\frac{1}{\alpha_{\texttt{RORO}}}\right)=\frac{p_{\max}-p_{\min}-2\gamma}{\nicefrac{{p_{\max}}}{{\alpha_{\texttt{RORO}}}}-p_{\max}-2\gamma}$ [Lechowicz:24].
@@
Suppose the $j^{\text{th}}$ instance of RORO has size $S^{(j)}=2^{j}$. From the start of the $j^{\text{th}}$ “round” of descending prices, the adversary begins presenting flexible demands with size smaller than or equal to $\sigma$ and a uniform deadline of $T$, until the total size of flexible demands presented so far is equal to $S^{(j)}-\sigma+\upsilon$, for small $\upsilon>0$. Note that at this point, the total size of flexible demands assigned to $\texttt{instance}_{j}$ is exactly $S^{(j)}-\sigma+\upsilon$. For sufficiently small $\iota$, the adversary reaches this point before the descending prices reach $p_{\min}$ – for the rest of the “round”, the adversary presents zero new flexible demand.
@@
Lemma D.1 ([Lechowicz:24, Lemma B.2]).
@@
Lemma D.2 ([Lechowicz:24, Lemma B.3]).
@@
Lemma D.1 suggests that we should be able to lower bound $\texttt{OPT}^{(j)}$ by $D^{(j)}\left[\phi^{(j)}(D^{(j)})-\gamma\right]$, where $D^{(j)}$ is the total amount of demand satisfied by $\texttt{instance}_{j}$. However, this is not true in OSDM– we already established that $\texttt{OPT}^{(j)}=D^{(j)}p_{\min}$, and the definition of the pseudo-cost function $\phi^{(j)}$ ensures that $\phi^{(j)}(S^{(j)})-\gamma=p_{\min}$. However, since $D^{(j)}<S^{(j)}$ (because $D^{(j)}=S^{(j)}-\sigma+\upsilon$) and $\phi^{(j)}$ is decreasing, we have $\phi^{(j)}(D^{(j)})-\gamma>p_{\min}$, so the bound does not hold. Thus, we must modify the lower bound on $\texttt{OPT}^{(j)}$ to account for the difference between $D^{(j)}$ and $S^{(j)}$.
@@
$\texttt{OPT}^{(j)}=D^{(j)}\left[\phi^{(j)}(\bar{w}^{(j)}+(S^{(j)}-D^{(j)}))-\gamma\right].$ (10)
@@
In the case of unit-size demand, it is given that $\int_{0}^{w}\phi(u)du+\gamma w+(1-w)p_{\max}=\alpha_{\texttt{RORO}}(\phi(w)-\gamma)$ for all $w\in[0,1]$ [Lechowicz:24, Theorem 3.2]. Extending this to the case of instance $j$ with size $S^{(j)}$, we have the following:
@@
$\texttt{RORO}^{(j)}=\alpha_{\texttt{RORO}}\frac{S^{(j)}}{D^{(j)}}\texttt{OPT}^{(j)}+\alpha_{\texttt{RORO}}S^{(j)}\left(p_{\max}-2\gamma-\frac{p_{\max}}{\alpha_{\texttt{RORO}}}\right)\exp\left(\frac{S^{(j)}-D^{(j)}}{\alpha_{\texttt{RORO}}}\right)-(S^{(j)}-D^{(j)})p_{\max}.$ (11)
@@
We note that in a regime where $\sigma$ approaches $1$, the above lower bound on $\zeta$ is ill-defined. In the OCS problem, it is known that the worst-possible competitive ratio is $\frac{p_{\max}+2\gamma}{p_{\min}}$ [Lechowicz:24], which is achieved by any feasible algorithm that purchases at any time step (i.e., possibly incurring the worst cost $p_{\max}$), and incurs the worst-possible switching cost to do so (i.e., $2\gamma$).
In the OSDM problem with no delivery costs (i.e., $c=\varepsilon=0$), no storage, and only flexible demands, we have a similar result – any algorithm is $\frac{p_{\max}+2\gamma}{p_{\min}}$-competitive as long as it is feasible. Thus, we have the following breakpoint for the lower bound on $\zeta$:
@@ Appendix E Deferred Proofs from § 4 (Analysis of PAAD Algorithm)
In this section, we provide full proofs for the results in § 4, which describes and analyzes the PAAD algorithm for the OSDM problem. We start by providing deferred plots of the competitive ratio $\alpha$ as a function of the problem’s parameters to showcase the important dependencies in Fig. 12. We then prove Theorem 4.1, which states the competitive bound of PAAD for OSDM-S, before proving 4.3, which states the corresponding bound for OSDM-T.
(a) Price ratio $\tfrac{p_{\max}}{p_{\min}}$
@@
Figure 12: Plotting the competitive ratio $\alpha$ (defined in (2)) as a function of the parameter dependencies described in each subfigure. Unless a parameter is on an $x$-axis, we set $p_{\min}=10$, $p_{\max}=200$, $\gamma=5$, $\delta=1$, $c=0.2$, $\varepsilon=0.05$, and $T=48$.
E.1 Proof of Theorem 4.1
@@
$\texttt{PAAD}(\mathcal{I})$ $\leq\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\int_{0}^{\hat{w}_{b}^{(i,j)}}\kern-10.00002pt\phi_{b}^{(i,j)}(u)du+\sum_{\tau=1}^{T}\left(\int_{0}^{\hat{w}_{f}^{(\tau)}}\kern-10.00002pt\phi_{f}^{(\tau)}(u)du+\int_{0}^{\hat{v}_{f}^{(\tau)}}\kern-10.00002pt\psi^{(\tau)}(z)dz\right)$ (12)
$\quad+\left(D_{b}+D_{f}-\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}-\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)(p_{\max}+2\gamma)$ (13)
$\quad+D_{b}\left(\bar{p}(c+\varepsilon)+2\delta\right)+\left(D_{f}-\sum_{\tau=1}^{T}\hat{v}_{f}^{(\tau)}\right)(p_{\max}(c+\varepsilon)+2\delta)$ (14)
$\quad-cp_{\min}\left(\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}+\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)+\hat{s}p_{\max}.$ (15)
@@
Using the previous lemmas, we now prove the competitive ratio. First, in the simple case where $D_{b}+D_{f}=0$, we have that the optimal cost is 0 and the cost of PAAD is at most $\hat{s}p_{\max}\leq Sp_{\max}$, so PAAD is $\alpha$-competitive under Definition 2.1 by setting $C=Sp_{\max}$. We thus focus on the more general case in which $D_{b}+D_{f}>0$. Using the results in Lemmas E.3, E.4, 4.4, and 4.5, we claim that the following holds:
@@
$\frac{Q+({\bm{\beta}}+D_{f}-\hat{W})(p_{\max}+2\gamma)+{\bm{\beta}}\left(\bar{p}(c+\varepsilon)+2\delta\right)+(D_{f}-\hat{V}_{f})(p_{\max}(c+\varepsilon)+2\delta)-cp_{\min}\hat{W}}{\sum_{i=1}^{n}G_{i}(\beta_{i})+\varepsilon\bar{p}{\bm{\beta}}+\sum_{\tau=1}^{T}(1+\varepsilon)H_{\tau}(f_{\tau})+(D_{f}+{\bm{\beta}})\frac{2\kappa}{T}}>\alpha.$ (16)
@@
Given the result in Lemma E.5, we proceed to work with the expression in terms of $D_{b}$. During the $i^{\text{th}}$ active period and the lifetime of the $j^{\text{th}}$ base demand driver, the minimum marginal purchasing price observed is given by $\phi_{b}^{(i,j)}(\hat{w}_{b}^{(i,j)}-2\gamma$ by the definition of the threshold function. Similarly, during the lifetime of the $\tau^{\text{th}}$ flexible demand driver, the minimum marginal purchasing and delivery cost observed is given by $\frac{1+\varepsilon}{1+c+\varepsilon}\left(\phi_{f}^{(\tau)}(\hat{w}_{f}^{(\tau)})+\psi_{f}^{(\tau)}(\hat{v}_{f}^{(\tau)})-2\kappa\right)$.
@@
$=\frac{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[X_{b}^{(i,j)}\right]+\sum_{\tau=1}^{T}\left[X_{f}^{(\tau)}\right]}{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[\left(\phi_{b}^{(i,j)}(\hat{w}_{b}^{(i,j)})\!-\!2\gamma\!+\!\varepsilon\bar{p}\!+\!\frac{2\kappa}{T}\right)B_{b}^{(i,j)}\right]\!+\!\sum_{\tau=1}^{T}\left[\frac{1+\varepsilon}{1+c+\varepsilon}\left(\phi_{f}^{(\tau)}(\hat{w}_{f}^{(\tau)})\!+\!\psi_{f}^{(\tau)}(\hat{v}_{f}^{(\tau)})\!-\!2\kappa\right)f_{\tau}\!+\!f_{\tau}\frac{2\kappa}{T}\right]},$ (17)
@@
But, if either of these two cases are true (i.e., for any unit of base or flexible demand), they contradict Lemmas 4.4 and 4.5, respectively. Thus, we have a contradiction, and the original assumption that (17) is $\alpha$ must be false. This completes the proof that $\frac{\texttt{PAAD}(\mathcal{I})-p_{\max}\hat{s}}{\texttt{OPT}(\mathcal{I})}\leq\alpha$.
@@
$\alpha_{\texttt{T}}\geq\frac{\omega}{\left[W\left(-\frac{\left((1+c+\varepsilon)p_{\max}-(1+\varepsilon)p_{\min}\right)e^{\frac{\omega\nicefrac{{2\delta}}{{T}}-p_{\min}c-(1+c+\varepsilon)p_{\max}}{(1+c+\varepsilon)p_{\max}+2(\eta+\delta)}}}{(1+c+\varepsilon)p_{\max}+2(\eta+\delta)}\right)+\frac{(1+c+\varepsilon)p_{\max}+p_{\min}c-\omega\nicefrac{{2\delta}}{{T}}}{(1+c+\varepsilon)p_{\max}+2(\eta+\delta)}\right]},$ (18)
@@
$\texttt{PAAD}(\mathcal{I})$ $\leq\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\int_{0}^{\hat{w}_{b}^{(i,j)}}\kern-10.00002pt\phi_{b}^{(i,j)}(u)du+\sum_{\tau=1}^{T}\left(\int_{0}^{\hat{w}_{f}^{(\tau)}}\kern-10.00002pt\phi_{f}^{(\tau)}(u)du+\int_{0}^{\hat{v}_{f}^{(\tau)}}\kern-10.00002pt\psi^{(\tau)}(z)dz\right)$ (19)
$\quad+\left(D_{b}+D_{f}-\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}-\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)(p_{\max}+2\eta)$ (20)
$\quad+D_{b}\left(\bar{p}(c+\varepsilon)+2\delta\right)+\left(D_{f}-\sum_{\tau=1}^{T}\hat{v}_{f}^{(\tau)}\right)(p_{\max}(c+\varepsilon)+2\delta)$ (21)
$\quad-cp_{\min}\left(\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}+\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)+\hat{s}p_{\max}.$ (22)
@@
The proof follows by the same logic as Lemma E.4 – the main change is with respect to (20), to capture the worst-case purchasing cost incurred by PAAD for purchasing the remaining demand that has not been purchased by the drivers.
@@
$\frac{Q+({\bm{\beta}}+D_{f}-\hat{W})(p_{\max}+2\eta)+{\bm{\beta}}\left(\bar{p}(c+\varepsilon)+2\delta\right)+(D_{f}-\hat{V}_{f})(p_{\max}(c+\varepsilon)+2\delta)-cp_{\min}\hat{W}}{\sum_{i=1}^{n}G_{i}(\beta_{i})+\varepsilon\bar{p}{\bm{\beta}}+\sum_{\tau=1}^{T}(1+\varepsilon)H_{\tau}(f_{\tau})+(D_{f}+{\bm{\beta}})\frac{2\delta}{T}}>\alpha_{\texttt{T}}.$ (23)
@@
$=\frac{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[X_{b}^{(i,j)}\right]+\sum_{\tau=1}^{T}\left[X_{f}^{(\tau)}\right]}{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[\left(\phi_{b}^{(i,j)}(\hat{w}_{b}^{(i,j)})\!-\!2\eta\!+\!\varepsilon\bar{p}\!+\!\frac{2\delta}{T}\right)B_{b}^{(i,j)}\right]\!+\!\sum_{\tau=1}^{T}\left[\frac{1+\varepsilon}{1+c+\varepsilon}\left(\phi_{f}^{(\tau)}(\hat{w}_{f}^{(\tau)})\!+\!\psi_{f}^{(\tau)}(\hat{v}_{f}^{(\tau)})\!-\!2(\eta\!+\!\delta)\right)f_{\tau}\!+\!f_{\tau}\frac{2\delta}{T}\right]},$ (24)

### v2

@@ Appendix D Deferred Proofs from § 3
In the above, $\alpha_{\texttt{RORO}}$ is the optimal competitive ratio for OCS, which is defined in the context of OSDM as the unique positive solution to $\exp\left(\frac{1}{\alpha_{\texttt{RORO}}}\right)=\frac{p_{\max}-p_{\min}-2\gamma}{\nicefrac{{p_{\max}}}{{\alpha_{\texttt{RORO}}}}-p_{\max}-2\gamma}$ (Lechowicz et al., 2024).
@@
Suppose the $j^{\text{th}}$ instance of RORO has size $S^{(j)}=2^{j}$. From the start of the $j^{\text{th}}$ “round” of descending prices, the adversary begins presenting flexible demands with size smaller than or equal to $\sigma$ and a uniform deadline of $T$, until the total size of flexible demands presented so far is equal to $S^{(j)}-\sigma+\upsilon$, for small $\upsilon>0$. Note that at this point, the total size of flexible demands assigned to $\texttt{instance}_{j}$ is exactly $S^{(j)}-\sigma+\upsilon$. For sufficiently small $\iota$, the adversary reaches this point before the descending prices reach $p_{\min}$—for the rest of the “round”, the adversary presents zero new flexible demand.
@@
Lemma D.1 ((Lechowicz et al., 2024, Lemma B.2)).
@@
Lemma D.2 ((Lechowicz et al., 2024, Lemma B.3)).
@@
Lemma D.1 suggests that we should be able to lower bound $\texttt{OPT}^{(j)}$ by $D^{(j)}\left[\phi^{(j)}(D^{(j)})-\gamma\right]$, where $D^{(j)}$ is the total amount of demand satisfied by $\texttt{instance}_{j}$. However, this is not true in OSDM—we already established that $\texttt{OPT}^{(j)}=D^{(j)}p_{\min}$, and the definition of the pseudo-cost function $\phi^{(j)}$ ensures that $\phi^{(j)}(S^{(j)})-\gamma=p_{\min}$. However, since $D^{(j)}<S^{(j)}$ (because $D^{(j)}=S^{(j)}-\sigma+\upsilon$) and $\phi^{(j)}$ is decreasing, we have $\phi^{(j)}(D^{(j)})-\gamma>p_{\min}$, so the bound does not hold. Thus, we must modify the lower bound on $\texttt{OPT}^{(j)}$ to account for the difference between $D^{(j)}$ and $S^{(j)}$.
@@
(11) $\texttt{OPT}^{(j)}=D^{(j)}\left[\phi^{(j)}(\bar{w}^{(j)}+(S^{(j)}-D^{(j)}))-\gamma\right].$
@@
In the case of unit-size demand, it is given that $\int_{0}^{w}\phi(u)du+\gamma w+(1-w)p_{\max}=\alpha_{\texttt{RORO}}(\phi(w)-\gamma)$ for all $w\in[0,1]$ (Lechowicz et al., 2024, Theorem 3.2). Extending this to the case of instance $j$ with size $S^{(j)}$, we have the following:
@@
(12) $\texttt{RORO}^{(j)}=\alpha_{\texttt{RORO}}\frac{S^{(j)}}{D^{(j)}}\texttt{OPT}^{(j)}+\alpha_{\texttt{RORO}}S^{(j)}\left(p_{\max}-2\gamma-\frac{p_{\max}}{\alpha_{\texttt{RORO}}}\right)\exp\left(\frac{S^{(j)}-D^{(j)}}{\alpha_{\texttt{RORO}}}\right)-(S^{(j)}-D^{(j)})p_{\max}.$
@@
We note that in a regime where $\sigma$ approaches $1$, the above lower bound on $\zeta$ is ill-defined. In the OCS problem, it is known that the worst-possible competitive ratio is $\frac{p_{\max}+2\gamma}{p_{\min}}$ (Lechowicz et al., 2024), which is achieved by any feasible algorithm that purchases at any time step (i.e., possibly incurring the worst cost $p_{\max}$), and incurs the worst-possible switching cost to do so (i.e., $2\gamma$).
In the OSDM problem with no delivery costs (i.e., $c=\varepsilon=0$), no storage, and only flexible demands, we have a similar result—any algorithm is $\frac{p_{\max}+2\gamma}{p_{\min}}$-competitive as long as it is feasible. Thus, we have the following breakpoint for the lower bound on $\zeta$:
@@ Appendix E Deferred Proofs from § 4 (Analysis of PAAD Algorithm)
In this section, we provide full proofs for the results in § 4, which describes and analyzes the PAAD algorithm for the OSDM problem. We start by providing deferred plots of the competitive ratio $\alpha$ as a function of the problem’s parameters to showcase the important dependencies in Fig. 12. We then prove Theorem 4.3, which states the competitive bound of PAAD for OSDM-S, before proving 4.4, which states the corresponding bound for OSDM-T.
(a) Price bound ratio $\tfrac{p_{\min}}{p_{\max}}$
@@
Figure 12. Plotting the competitive ratio $\alpha$ (defined in (3)) as a function of the parameter dependencies described in each subfigure. Unless a parameter is on an $x$-axis, we set $p_{\min}=10$, $p_{\max}=200$, $\gamma=5$, $\delta=1$, $c=0.2$, $\varepsilon=0.05$, and $T=48$.
E.1. Proof of Theorem 4.3
@@
(13) $\texttt{PAAD}(\mathcal{I})$ $\leq\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\int_{0}^{\hat{w}_{b}^{(i,j)}}\kern-10.00002pt\phi_{b}^{(i,j)}(u)du+\sum_{\tau=1}^{T}\left(\int_{0}^{\hat{w}_{f}^{(\tau)}}\kern-10.00002pt\phi_{f}^{(\tau)}(u)du+\int_{0}^{\hat{v}_{f}^{(\tau)}}\kern-10.00002pt\psi^{(\tau)}(z)dz\right)$
(14) $\quad+\left(D_{b}+D_{f}-\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}-\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)(p_{\max}+2\gamma)$
(15) $\quad+D_{b}\left(\bar{p}(c+\varepsilon)+2\delta\right)+\left(D_{f}-\sum_{\tau=1}^{T}\hat{v}_{f}^{(\tau)}\right)(p_{\max}(c+\varepsilon)+2\delta)$
(16) $\quad-cp_{\min}\left(\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}+\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)+\hat{s}p_{\max}.$
@@
Using the previous lemmas, we now prove the competitive ratio. First, in the simple case where $D_{b}+D_{f}=0$, we have that the optimal cost is 0 and the cost of PAAD is at most $\hat{s}p_{\max}\leq Sp_{\max}$, so PAAD is $\alpha$-competitive under Definition 2.1 by setting $C=Sp_{\max}$. We thus focus on the more general case in which $D_{b}+D_{f}>0$. Using the results in Lemmas E.3, E.4, 4.5, and 4.6, we claim that the following holds:
@@
Further, let $\hat{V}_{f}=\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{v}_{f}^{(\tau)}$ denote the total delivery by all flexible demand drivers.
@@
(17) $\frac{Q+({\bm{\beta}}+D_{f}-\hat{W})(p_{\max}+2\gamma)+{\bm{\beta}}\left(\bar{p}(c+\varepsilon)+2\delta\right)+(D_{f}-\hat{V}_{f})(p_{\max}(c+\varepsilon)+2\delta)-cp_{\min}\hat{W}}{\sum_{i=1}^{n}G_{i}(\beta_{i})+\varepsilon\bar{p}{\bm{\beta}}+\sum_{\tau=1}^{T}(1+\varepsilon)H_{\tau}(f_{\tau})+(D_{f}+{\bm{\beta}})\frac{2\kappa}{T}}>\alpha.$
@@
Given the result in Lemma E.5, we proceed to work with the expression in terms of $D_{b}$. During the $i^{\text{th}}$ active period and the lifetime of the $j^{\text{th}}$ base demand driver, the minimum marginal purchasing price observed is given by $\phi_{b}^{(i,j)}(\hat{w}_{b}^{(i,j)})-2\gamma$ by the definition of the threshold function. Similarly, during the lifetime of the $\tau^{\text{th}}$ flexible demand driver, the minimum marginal purchasing and delivery cost observed is given by $\frac{1+\varepsilon}{1+c+\varepsilon}\left(\phi_{f}^{(\tau)}(\hat{w}_{f}^{(\tau)})+\psi_{f}^{(\tau)}(\hat{v}_{f}^{(\tau)})-2\kappa\right)$.
@@
(18) $=\frac{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[X_{b}^{(i,j)}\right]+\sum_{\tau=1}^{T}\left[X_{f}^{(\tau)}\right]}{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[\left(\phi_{b}^{(i,j)}(\hat{w}_{b}^{(i,j)})\!-\!2\gamma\!+\!\varepsilon\bar{p}\!+\!\frac{2\kappa}{T}\right)B_{b}^{(i,j)}\right]\!+\!\sum_{\tau=1}^{T}\left[\frac{1+\varepsilon}{1+c+\varepsilon}\left(\phi_{f}^{(\tau)}(\hat{w}_{f}^{(\tau)})\!+\!\psi_{f}^{(\tau)}(\hat{v}_{f}^{(\tau)})\!-\!2\kappa\right)f_{\tau}\!+\!f_{\tau}\frac{2\kappa}{T}\right]},$
@@
But, if either of these two cases are true (i.e., for any unit of base or flexible demand), they contradict Lemmas 4.5 and 4.6, respectively. Thus, we have a contradiction, and the original assumption that (18) is $\alpha$ must be false. This completes the proof that $\frac{\texttt{PAAD}(\mathcal{I})-p_{\max}\hat{s}}{\texttt{OPT}(\mathcal{I})}\leq\alpha$.
@@
(19) $\alpha_{\texttt{T}}\geq\frac{\omega}{\left[W\left(-\frac{\left((1+c+\varepsilon)p_{\max}-(1+\varepsilon)p_{\min}\right)e^{\frac{\nicefrac{{\omega 2\delta}}{{T}}-p_{\min}c-(1+c+\varepsilon)p_{\max}}{(1+c+\varepsilon)p_{\max}+2(\eta+\delta)}}}{(1+c+\varepsilon)p_{\max}+2(\eta+\delta)}\right)+\frac{(1+c+\varepsilon)p_{\max}+p_{\min}c-\nicefrac{{\omega 2\delta}}{{T}}}{(1+c+\varepsilon)p_{\max}+2(\eta+\delta)}\right]},$
@@
(20) $\texttt{PAAD}(\mathcal{I})$ $\leq\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\int_{0}^{\hat{w}_{b}^{(i,j)}}\kern-10.00002pt\phi_{b}^{(i,j)}(u)du+\sum_{\tau=1}^{T}\left(\int_{0}^{\hat{w}_{f}^{(\tau)}}\kern-10.00002pt\phi_{f}^{(\tau)}(u)du+\int_{0}^{\hat{v}_{f}^{(\tau)}}\kern-10.00002pt\psi^{(\tau)}(z)dz\right)$
(21) $\quad+\left(D_{b}+D_{f}-\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}-\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)(p_{\max}+2\eta)$
(22) $\quad+D_{b}\left(\bar{p}(c+\varepsilon)+2\delta\right)+\left(D_{f}-\sum_{\tau=1}^{T}\hat{v}_{f}^{(\tau)}\right)(p_{\max}(c+\varepsilon)+2\delta)$
(23) $\quad-cp_{\min}\left(\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}+\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)+\hat{s}p_{\max}.$
@@
The proof follows by the same logic as Lemma E.4—the main change is with respect to (21), to capture the worst-case purchasing cost incurred by PAAD for purchasing the remaining demand that has not been purchased by the drivers.
@@
(24) $\frac{Q+({\bm{\beta}}+D_{f}-\hat{W})(p_{\max}+2\eta)+{\bm{\beta}}\left(\bar{p}(c+\varepsilon)+2\delta\right)+(D_{f}-\hat{V}_{f})(p_{\max}(c+\varepsilon)+2\delta)-cp_{\min}\hat{W}}{\sum_{i=1}^{n}G_{i}(\beta_{i})+\varepsilon\bar{p}{\bm{\beta}}+\sum_{\tau=1}^{T}(1+\varepsilon)H_{\tau}(f_{\tau})+(D_{f}+{\bm{\beta}})\frac{2\delta}{T}}>\alpha_{\texttt{T}}.$
@@
(25) $=\frac{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[X_{b}^{(i,j)}\right]+\sum_{\tau=1}^{T}\left[X_{f}^{(\tau)}\right]}{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[\left(\phi_{b}^{(i,j)}(\hat{w}_{b}^{(i,j)})\!-\!2\eta\!+\!\varepsilon\bar{p}\!+\!\frac{2\delta}{T}\right)B_{b}^{(i,j)}\right]\!+\!\sum_{\tau=1}^{T}\left[\frac{1+\varepsilon}{1+c+\varepsilon}\left(\phi_{f}^{(\tau)}(\hat{w}_{f}^{(\tau)})\!+\!\psi_{f}^{(\tau)}(\hat{v}_{f}^{(\tau)})\!-\!2(\eta\!+\!\delta)\right)f_{\tau}\!+\!f_{\tau}\frac{2\delta}{T}\right]},$